### v3

# Graph Learning Across Data Silos

Abstract
We consider the problem of inferring graph topology from smooth graph signals in a novel but practical scenario where data are located in distributed clients and prohibited from leaving local clients due to factors such as privacy concerns. The main difficulty in this task is how to exploit the potentially heterogeneous data of all clients under data silos. To this end, we first propose an auto-weighted multiple graph learning model to jointly learn a personalized graph for each local client and a single consensus graph for all clients. The personalized graphs match local data distributions, thereby mitigating data heterogeneity, while the consensus graph captures the global information. Moreover, the model can automatically assign appropriate contribution weights to local graphs based on their similarity to the consensus graph. We next devise a tailored algorithm to solve the induced problem, where all raw data are processed locally without leaving clients. Theoretically, we establish a provable estimation error bound and convergence analysis for the proposed model and algorithm. Finally, extensive experiments on synthetic and real data are carried out, and the results illustrate that our approach can learn graphs effectively in the target scenario.
Index Terms:
Graph topology inference, multiple graph learning, graph signal processing, data silos, privacy-preserving
I Introduction
Graphs are powerful tools for flexibly describing topological relationships between data entities [1], which have been extensively applied in many celebrated models, e.g., spectral clustering [2] and graph neural networks [3]. Among these applications, a graph that accurately represents the information inherent in the structured data is required, which, however, is not available in many cases. An alternative approach is to learn graphs directly from raw data, termed graph learning (GL), for downstream tasks [4, 1]. Recently, with the rise of graph signal processing (GSP) [5], various methods attempted to learn graphs from a signal processing perspective. One of the most studied GSP-based models—the smoothness-based model—postulates that the observed graph signals are smooth over the graph to be learned [6, 7]. Intuitively, a smooth graph signal means that the signal values corresponding to two connected nodes of the underlying graph are similar[7]. Typically, learning graphs from smooth signals is equivalent to minimizing a constrained quadratic Laplacian form problem, where the quadratic Laplacian form is related to signal smoothness, and the constraints (regularizers) are used to assign properties to the graphs, such as sparsity [6] and node connectivity [7]. Many signals appear to be smooth over their underlying graphs, e.g., meteorology data [6] and medical data [8], implying numerous applications of smoothness-based GL.
Here, we consider learning graphs in a previously unexplored scenario, where data are stored across different clients (e.g., companies and organizations) and are not allowed to leave their clients due to factors such as privacy concerns. This scenario is known as data-silos [9], and a typical example is medical data. Suppose some hospitals collect brain fMRI data from autistic and non-autistic individuals separately to learn the impact of autism on connectivity networks of brain functional regions [8]. However, the data are prohibited from leaving the hospital where they are stored as people are reluctant to disclose their private data. A naive way is for each client to infer graph topology independently, which avoids any data leakage. Although feasible, this approach ignores potential topological relationships between local graphs, resulting in suboptimal results. Thus, it is more reasonable to learn graphs using data information from all clients. The task is not trivial and faces two main challenges. The first challenge is how to leverage siloed data from all clients collaboratively. Furthermore, graph signals across different silos are inherently non-IID due to factors such as different data collection protocols and heterogeneous underlying graphs [10]. We still use the brain fMRI data as an example. The brain functional connectivity networks of autistic and non-autistic individuals are different due to the impact of autism. It is unreasonable to utilize heterogeneous data to learn a single global graph like traditional paradigms [6, 7]. Thus, the second challenge is handling heterogeneous data.
Regarding the first challenge, federated learning (FL)[11, 12, 13] is an emerging tool for learning models based on datasets distributed across multiple clients under privacy constraints. The primary feature of FL is that clients transmit model updates instead of raw data to a central server to collaboratively learn a global model [11]. Privacy can be preserved in this schema to some extent since all data are processed locally [12]. However, traditional FL algorithms, e.g., FedAvg [14], learn a single global model from all data, which may suffer from performance degradation when data across different clients are heterogeneous. A widely used approach to handle data heterogeneity—the second challenge—is personalized FL (PFL) [15]. The philosophy behind PFL is that we learn for each client a personalized model that matches its data distribution to mitigate the impact of heterogeneity. The techniques for adapting global models for individual clients include transfer learning [16], multi-task learning [17], and meta-learning [18].
Borrowing the idea from PFL, we learn a personalized graph for each local client to handle data heterogeneity. However, it poses two additional challenges due to the characteristics of GL tasks. (i) Our goal is to learn all local graphs jointly by exploiting their latent relationships so that each local graph benefits from ”borrowing” information from other datasets. Thus, it is crucial to describe the relationships between local graphs, which is the focus of multiple graph learning (MGL). A common approach is to design regularization penalties, e.g., fused Lasso penalty [19], group Lasso penalty [20], Gram matrix-based penalty [21], and those describing temporally topological relationships [22, 23, 24], etc. These regularizers characterize topological relationships among multiple graphs, but few consider capturing common structures from all local graphs. In practice, many graphs share common structures. For example, connections between normally functioning brain regions in autistic and non-autistic individuals should remain the same. The common structures reflect the global information across all local datasets, which may be useful for many downstream tasks. (ii) It is infeasible to apply existing PFL algorithms, such as [25, 26, 17, 27, 28], to our task since their personalization methods do not fit the GL problem. Besides, the problem of concern can be categorized as cross-silo FL, where clients are a few companies or organizations rather than massive devices in cross-device FL [11, 14]. Thus, a tailored algorithm is required to learn graphs in the target scenario.
To address these issues, we propose a framework to learn graphs from smooth but heterogeneous data under data silos. Our contributions can be summarized as follows:
  1. $\bullet$
    We propose an auto-weighted MGL model in which local personalized graphs and a consensus graph are jointly learned. The consensus graph can capture common structures representing global information across all datasets, while the local graphs preserve the heterogeneity of local datasets. Furthermore, our model can automatically assign contribution weights to local graphs based on their similarity to the consensus graph. Theoretically, we provide the estimation error bound of the proposed method to reveal some key factors affecting graph estimation performance.
  2. $\bullet$
    We develop a privacy-preserving algorithm to learn graphs in the target scenario. Our algorithm follows the communication protocol of FL, where model updates instead of raw data are transmitted to a central server to learn all graphs collaboratively. A provable convergence analysis of the algorithm is provided, illustrating that the convergence rate of our algorithms is $\mathcal{O}(1/T)$, where $T$ is the number of communication rounds.
  3. $\bullet$
    Extensive experiments with synthetic and real-world data are conducted to validate our framework, and the results show that our approach can effectively learn graphs under data silos.
Organization: The rest of this paper is organized as follows. We start with background information and the problem of concern in Section II. The proposed model for learning graphs in the target scenario is presented in III, followed by the corresponding algorithm in Section IV. Experimental setups and results are provided in Section V. Finally, concluding remarks are presented in Section VI.
[…]
II Background and Problem Statement
II-A GSP Background
We consider undirected graphs with non-negative weights and no self-loops. For such a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ with $d$ vertices, where $\mathcal{V}$ and $\mathcal{E}$ are the sets of vertices and edges, respectively, its adjacency matrix $\mathbf{A}\in\mathbb{S}^{d\times d}$ is a symmetric matrix with zero diagonal entries and non-negative off-diagonal entries. The Laplacian matrix of $\mathcal{G}$ is $\mathbf{L}=\mathbf{D}-\mathbf{A}$ [29], where the degree matrix $\mathbf{D}\in\mathbb{S}^{d\times d}$ is a diagonal matrix satisfying $\mathbf{D}{[ii]}=\sum_{j=1}^{d}\mathbf{A}{[ij]}$. The matrices $\mathbf{A}$ and $\mathbf{L}$ encode the topology of $\mathcal{G}$ since they have a one-to-one relationship. We study the graph signal $\mathbf{x}=\left[\mathbf{x}{[1]},...,\mathbf{x}{[d]}\right]^{\top}\in\mathbb{R}^{d}$ associated with $\mathcal{G}$, where $\mathbf{x}{[i]}$ is the signal value of node $i\in\mathcal{V}$. The smoothness of $\mathbf{x}$ over $\mathcal{G}$ is defined as follows.
Definition 1.
(Smoothness [6]). Given a graph signal $\mathbf{x}$ and a graph $\mathcal{G}$ whose Laplacian matrix and adjacency matrix are $\mathbf{L}$ and $\mathbf{A}$, respectively, the smoothness of $\mathbf{x}$ over $\mathcal{G}$ is
$\mathbf{x}^{\top}\mathbf{L}\mathbf{x}=\frac{1}{2}\sum_{i,j}\mathbf{A}{[ij]}\left(\mathbf{x}{[i]}-\mathbf{x}{[j]}\right)^{2}.$ (1)
The Laplacian quadratic form (1) is known as the Dirichlet energy, which quantifies how much the signal $\mathbf{x}$ changes w.r.t. $\mathcal{G}$. A small value of (1) indicates limited signal variability, meaning that $\mathbf{x}$ is smooth over the corresponding graph [6].
II-B Graph Learning From Smooth Signals
Given $N$ observations $\mathbf{X}=\left[\mathbf{x}_{1},...,\mathbf{x}_{N}\right]\in\mathbb{R}^{d\times N}$, smoothness-based GL aims to infer the underlying graph topology $\mathcal{G}$ under the assumption that the signals $\mathbf{X}$ are smooth over $\mathcal{G}$. Formally, the problem is written as
$\underset{\mathbf{L}\in\mathcal{L}}{\mathrm{min}}\,\,\frac{1}{N}\sum_{n=1}^{N}\mathbf{x}_{n}^{\top}\mathbf{L}\mathbf{x}_{n}-\alpha\mathbf{1}^{\top}\log\left(\mathrm{diag}(\mathbf{L})\right)+{\beta}\lVert\mathbf{L}\rVert^{2}_{\mathrm{F,off}},$ (2)
where the first term is to quantify the smoothness of $\mathbf{X}$ over $\mathcal{G}$, and the last two terms are regularizers that endow $\mathcal{G}$ with desired properties. The first and second terms of the regularizers control node connectivity and edge sparsity of the learned graph [7], where $\alpha$ and $\beta$ are predefined constants. We use Laplacian matrix $\mathbf{L}$ to represent graph topology of $\mathcal{G}$, which lies in the following set $\mathcal{L}$
$\mathcal{L}\triangleq\left\{\mathbf{L}:\mathbf{L}\in\mathbb{S}^{d\times d},\,\mathbf{L}\mathbf{1}=\mathbf{0},\,\mathbf{L}{[ij]}\leq 0\,\,\text{for}\,\,i\neq j\right\}.$ (3)
Based on (1), problem (2) can be rephrased as
$\underset{\mathbf{A}\in\mathcal{A}}{\mathrm{min}}\,\,\frac{1}{2N}\lVert\mathbf{A}\circ\mathbf{Z}\rVert_{1,1}-\alpha\mathbf{1}^{\top}\mathrm{log}(\mathbf{A}\mathbf{1})+{\beta}\lVert\mathbf{A}\rVert_{\mathrm{F}}^{2},$ (4)
where $\lVert\cdot\rVert_{1,1}$ is the element-wise $\ell_{1}$ norm of a matrix. Besides, $\mathbf{Z}\in\mathbb{R}^{d\times d}$ is a pairwise distance matrix defined as
$\mathbf{Z}{[ij]}=\lVert\widetilde{\mathbf{x}}_{i}-\widetilde{\mathbf{x}}_{j}\rVert_{2}^{2},$ (5)
where $\widetilde{\mathbf{x}}_{i}\in\mathbb{R}^{N}$ is the $i$-th row vector of $\mathbf{X}$. Similarly, $\mathcal{A}$ is the set containing all adjacency matrices,
$\mathcal{A}=\left\{\mathbf{A}:\mathbf{A}\in\mathbb{S}^{d\times d},\mathbf{A}\geq 0,\mathrm{diag}(\mathbf{A})=\mathbf{0}\right\},$ (6)
where $\mathbf{A}\geq 0$ means that all elements of $\mathbf{A}$ are non-negative. By the definition of $\mathcal{A}$, the number of free variables of $\mathbf{A}$ is $p:=\frac{d(d-1)}{2}$ [7]. For simplicity, we define a vector $\mathbf{w}\in\mathbb{R}^{p}$ whose elements are the upper triangle variables of $\mathbf{A}$. Then, problem (4) can be rewritten into a vector form as
$\underset{\mathbf{w}\geq 0}{\mathrm{min}}\,\,\frac{1}{N}\mathbf{z}^{\top}\mathbf{w}{-\alpha\mathbf{1}^{\top}\mathrm{log}(\mathbf{S}\mathbf{w})+2\beta\lVert\mathbf{w}\rVert_{2}^{2}},$ (7)
where $\mathbf{S}$ is a linear operator satisfying $\mathbf{S}\mathbf{w}=\mathbf{A}\mathbf{1}$, and $\mathbf{z}$ is the vector form of the upper triangle elements of $\mathbf{Z}$ [^1] The vector $\mathbf{z}$ is calculated from $\mathbf{x}_{1},...,\mathbf{x}_{N}$. Thus, $\mathbf{z}$ is also referred to as observation data in the following sections..
Figure 1: The illustration of the target scenario. The clients $\mathcal{C}_{1},...\mathcal{C}_{I}$ store graphs signals $\mathbf{X}_{1},...,\mathbf{X}_{I}$ generated from $I$ distinct but related graphs. The data are not allowed to leave their clients, but all clients can commute with a central server.
II-C Problem Statement
As shown in Fig.1, assuming there are $I$ clients $\mathcal{C}_{1},...,\mathcal{C}_{I}$, the $i$-th client stores signals $\mathbf{X}_{i}\in\mathbb{R}^{d\times N_{i}}$ generated from the graph $\mathcal{G}_{i}$, where $N_{i}$ is the size of the $i$-th dataset. All graphs are defined over the same node set, and the data are prohibited from leaving the client where they are stored due to factors such as privacy concerns. However, the clients can exchange information with a central server. We assume that (i) graph signals $\mathbf{X}_{i}$ are smooth over $\mathcal{G}_{i}$, and (ii) $\mathcal{G}_{1},...,\mathcal{G}_{I}$ may be heterogeneous, meaning that the corresponding graph signals could be non-IID. Our goal is to infer the graphs from $\mathbf{X}_{1},...,\mathbf{X}_{I}$ in the case of data silos.
III Model Formulation
In this section, we propose an auto-weighted MGL model to learn graphs in the target scenario.
III-A Basic Formulation
We assume that $I$ local graphs, which are denoted as $\mathbf{w}_{1},...,\mathbf{w}_{I}\in\mathbb{R}^{p}$, have some common structures named consensus graphs $\mathbf{w}_{\mathrm{con}}\in\mathbb{R}^{p}$. Intuitively, the consensus graph should be close to local graphs since they share common edges. Furthermore, the consensus graph should also be sparse to remove redundant noisy edges [30]. Thus, we learn $I$ local graphs and the consensus graph jointly via
$\underset{\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}\in\mathcal{W}}{\min}$ $\sum_{i=1}^{I}\underbrace{\frac{1}{N_{i}}\mathbf{z}_{i}^{\top}\mathbf{w}_{i}-\alpha\mathbf{1}^{\top}\log\left(\mathbf{S}\mathbf{w}_{i}\right)+2\beta\lVert\mathbf{w}_{i}\rVert_{2}^{2}}_{{g}_{i}(\mathbf{w}_{i})}$
$+{\lambda\nu\sum_{i=1}^{I}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}+\lambda\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}}$
$=\underset{\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}\in\mathcal{W}}{\min}$ $G(\mathbf{W})+\lambda R(\mathbf{W})$ (8)
where $\mathbf{z}_{i}$ is the observed data in client $\mathcal{C}_{i}$, $\mathcal{W}:=\left\{\mathbf{w}:\mathbf{w}\geq 0\right\}$, $\mathbf{W}=[\mathbf{w}_{1},...,\mathbf{w}_{I},\mathbf{w}_{\mathrm{con}}]\in\mathbb{R}^{p\times{(I+1)}}$, $G(\mathbf{W}):=\sum_{i=1}^{I}g_{i}(\mathbf{w}_{i})$, and $R(\mathbf{W}):=\nu\sum_{i=1}^{I}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}+\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}$. Our model consists of two parts. The first part $G(\mathbf{W})$ represents that local clients utilize the traditional smoothness-based model (7) to learn graphs $\mathbf{w}_{1},...,\mathbf{w}_{I}$ from locally stored data. The second term $R(\mathbf{W})$ is the regularizer that topologically connects local graphs via the consensus graph. Specifically, the first term $\nu\sum_{i=1}^{I}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}$ of $R(\mathbf{W})$ measures difference between local graphs and the consensus graph. We add the $\ell_{1}$ norm term because $\mathbf{w}_{\mathrm{con}}$ is expected to be sparse. The constants $\nu$ and $\lambda$ are predefined parameters. The parameters of problem (8) are $\alpha,\beta,\nu$, and $\lambda$. However, as stated in [7], tuning the importance of the log-degree term w.r.t. the other graph terms has a scaling effect. We can fix $\alpha$ and search for the other parameters. Therefore, the free parameters of our model are $\beta$ $\nu$ and $\lambda$.
The proposed model (8) enjoys the following advantages. Firstly, our model provides a way to learn local personalized graphs $\mathbf{w}_{1},...,\mathbf{w}_{I}$ jointly. Compared to learning graphs independently, our model utilizes the information from all datasets, which could boost learning performance. Secondly, we learn a personalized graph for each client, which alleviates the bias of learning a single global graph using all data in the case of data heterogeneity. Lastly, unlike most PFL methods that only learn personalized local models [15], our model learns a consensus graph reflecting global information.
III-B Auto-weighted Multiple Graph Learning
At first glance, our model (8) treats $I$ local graphs equally since they share the same weight in $\nu\sum_{i=1}^{I}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}$. This is unreasonable because the similarity between the consensus graph and different local graphs could vary widely. Fortunately, we will show that our model (8) can implicitly assign appropriate weights to local graphs, which we termed contribution weights, based on their similarity to $\mathbf{w}_{\mathrm{con}}$ via the inverse distance weighting schema [31].
Specifically, if we solve (8) by alternately updating $\mathbf{w}_{1},...,\mathbf{w}_{I}$ and $\mathbf{w}_{\mathrm{con}}$—which is exactly how we solve it in Section IV—when we fix the consensus graph $\mathbf{w}_{\mathrm{con}}$, the sub-problem of updating $\mathbf{w}_{i},i=1,...,I,$ is
$\underset{\mathbf{w}_{i}\in\mathcal{W}}{\min}\;\;g_{i}(\mathbf{w}_{i})+\rho\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2},$ (9)
where we let $\rho=\lambda\nu$ for simplicity. We take the derivative of the Lagrange function of (9) and set it to zero, which yields
$\rho\widetilde{\gamma}_{i}\frac{\partial\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}^{2}}{\partial\mathbf{w}_{i}}+\frac{\partial g_{i}(\mathbf{w}_{i})}{\partial\mathbf{w}_{i}}-\frac{\partial\bm{\theta}_{i}^{\top}\mathbf{w}_{i}}{\partial\mathbf{w}_{i}}=\mathbf{0},$ (10)
where $\bm{\theta}_{i}\in\mathbb{R}^{p}$ is the Lagrange multiplier, and
$\widetilde{\gamma}_{i}=\frac{1}{2\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}}.$ (11)
Note that (11) depends on $\mathbf{w}_{i}$, meaning that $\widetilde{\gamma}_{i}$ and $\mathbf{w}_{i}$ are coupled with each other. However, if we set $\widetilde{\gamma}_{i}$ stationary, (10) is the solution to the following problem
$\underset{\mathbf{w}_{i}\in\mathcal{W}}{\min}\;\;g_{i}(\mathbf{w}_{i})+\rho\widetilde{\gamma}_{i}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}^{2}.$ (12)
After solving (12), we can use the obtained $\mathbf{w}_{i}$ to update the weight $\widetilde{\gamma}_{i}$ via (11). Therefore, we can alternately update $\mathbf{w}_{i}$ and $\widetilde{\gamma}_{i}$ to solve (9). Similarly, it is not difficult to check that we can update $\mathbf{w}_{\mathrm{con}}$ using the same strategy as updating $\mathbf{w}_{i}$.
Combing the alternative updates of $\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}$, and $\widetilde{\gamma}_{i}$, the basic formulation (8) is rephrased as
$\underset{\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}}{\min}$ $\sum_{i=1}^{I}{\frac{1}{N_{i}}\mathbf{z}_{i}^{\top}\mathbf{w}_{i}-\alpha\mathbf{1}^{\top}\log\left(\mathbf{S}\mathbf{w}_{i}\right)+2\beta\lVert\mathbf{w}_{i}\rVert_{2}^{2}}$
$+{\frac{\rho}{2}\sum_{i=1}^{I}{\gamma}_{i}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}^{2}+\lambda\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}}$
$\mathrm{s.t.}\;\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}\in\mathcal{W},\gamma_{i}=\frac{1}{\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}}.$ (13)
One merit of solving (8) via the reformulation (13) is that it naturally produces a contribution weight ${\gamma}_{i}$ for the local graph $\mathbf{w}_{i}$. The weight value is determined by the similarity between $\mathbf{w}_{i}$ and $\mathbf{w}_{\mathrm{con}}$. For the $\mathbf{w}_{i}$ close to $\mathbf{w}_{\mathrm{con}}$, a larger ${\gamma}_{i}$ is assigned to the corresponding term, increasing the contribution of the $i$-th local graph to the consensus graph. On the contrary, the local graph far from the consensus graph obtains a small ${\gamma}_{i}$. Thus, our model can implicitly and automatically assign appropriate weights to local clients based on their similarity to the consensus graph.
III-C Theoretical Analysis
Let $\mathbf{W}^{*}=[\mathbf{w}^{*}_{1},...,\mathbf{w}^{*}_{I},\mathbf{w}^{*}_{\mathrm{con}}]\in\mathbb{R}^{p\times{(I+1)}}$ be the true graphs and $\widehat{\mathbf{W}}=[\widehat{\mathbf{w}}_{1},...,\widehat{\mathbf{w}}_{I},\widehat{\mathbf{w}}_{\mathrm{con}}]\in\mathbb{R}^{p\times{(I+1)}}$ be the estimated graphs of our model. We aim to derive the estimation error bound of our proposed multiple-graph estimator, i.e., $\lVert\widehat{\mathbf{W}}-\mathbf{W}^{*}\rVert_{\mathrm{F}}$, by using the tools from [32]. Before conducting the analysis, we make the following assumptions.
Assumption 1.
The real covariance of signals in $\mathcal{C}_{i}$ is $\varpi(\mathbf{L}^{*}_{i})$, where $\mathbf{L}^{*}_{i}$ is the Laplacian matrix of $\mathbf{w}^{*}_{i}$, and $\sqrt{\varpi(\mathbf{L}^{*}_{i})}:\mathbb{R}^{d\times d}\to\mathbb{R}^{d\times d}$ is a graph filter defined on $\mathbf{L}^{*}_{i}$.
Assumption 2.
All real local graphs $\mathbf{w}^{*}_{1},...,\mathbf{w}^{*}_{I}$ have no isolated nodes, i.e., there exists a constant $B_{d}>0$ such that the node degrees of true graphs are greater than $B_{d}$. In addition, the real graphs have limited edge weights, i.e., there exists a constant $B_{w^{*}}>0$ such that $\lVert\mathbf{w}_{i}^{*}\rVert_{2}<B_{w^{*}},i=1,...,I$.
Assumption 1 is widely used to define smooth signals, and a typical $\varpi(\mathbf{L}^{*}_{i})$ is $(\mathbf{L}^{*}_{i})^{\dagger}$ [6, 7]. Assumption 2 holds when the true graphs have limited edge weights and no isolated nodes, which is common in the real world. Based on these assumptions, we provide the following theorem.
Theorem 1.
Let $s=\#\{(i,j):\mathbf{W}^{*}[ij]\neq 0,i=1,...,p,j=1,...,I+1\}$ denote the sparsity parameter, $N_{\min}=\underset{i=1,...,I}{\min}N_{i}$. For $N_{\min}\geq 2\ln d$ and $\lambda\geq 2\omega_{\min}(\mathbf{L}_{m})\sqrt{I+1}\left(C_{r}+160\sqrt{2}C_{\varpi}\sqrt{\frac{\ln d}{N_{\min}}}\right)$, we have probability at least $1-{4I}/{d}$ such that
$\lVert\widehat{\mathbf{W}}-\mathbf{W}^{*}\rVert_{\mathrm{F}}\leq$ $\frac{3\omega_{\min}(\mathbf{L}_{m})\sqrt{I+1}}{4\beta}\left(\sqrt{s}+\nu\sqrt{Ip\omega_{\max}(\mathbf{L}_{m})}\right)$
$\left(C_{r}+160\sqrt{2}C_{\varpi}\sqrt{\frac{\ln d}{N_{\min}}}\right),$ (14)
where
$\mathbf{L}_{m}\in\mathbb{R}^{(I+1)\times(I+1)}=\begin{bmatrix}1&0&\cdots&0&-1\\ 0&1&\cdots&0&-1\\ \vdots&\ddots&\cdots&0&-1\\ 0&0&\cdots&1&-1\\ -1&-1&-1&-1&I\end{bmatrix},$ (20)
$C_{r}=\sqrt{\frac{2d(d-1)}{B_{d}}}+4\beta B_{w^{*}}+4\max_{i}\left\lVert\varpi(\mathbf{L}_{i}^{*})\right\rVert_{\infty}$, $\left\lVert\varpi(\mathbf{L}_{i}^{*})\right\rVert_{\infty}$ returns the maximum absolute value of the elements in $\varpi(\mathbf{L}_{i}^{*})$, $C_{\varpi}=\max_{i,k}\varpi(\mathbf{L}_{i}^{*}){[kk]}$, $\omega_{\max}(\mathbf{L}_{m})$ and $\omega_{\min}(\mathbf{L}_{m})$ are respectively the maximum and minimum eigenvalues of $\mathbf{L}_{m}$.
Proof:
See Appendix A for details. ∎
Theorem 1 characterizes the estimation error bound of our proposed model w.r.t. some factors, such as graph size $d$(the number of the free variables $p$), the sparsity parameter $s$, the number of local graphs $I$, and some conditions on the real graphs. The theorem states that if we choose the regularizer weight $\lambda$ properly, the estimation error of our model is bounded by the r.h.s. of (14) with high probability. Existing MGL models, such as [21, 33], also derive their estimation error bounds using the tools in [32]. Our result and proof are different from those in [21, 33] since the basic GL model $G(\mathbf{W})$ and the regularizer $R(\mathbf{W})$ of our model have not been theoretically analyzed before.
III-D Connections to Existing MGL Models
In the literature, there have been some works that learn multiple graphs based on the assumption of common structures. These works differ mainly in how they describe structural relationships among multiple graphs. To avoid notational confusion, we use different notations to interpret existing works. We let $\mathbf{C}\in\mathbb{R}^{d\times d}$ denote the common graph structures and $\mathbf{U}_{1},...,\mathbf{U}_{I}\in\mathbb{R}^{d\times d}$ represent the unique structures of $I$ local graphs. Under these notations, the graph regularizer $R(\mathbf{W})$ in our model (8) can be roughly summarized as (not exactly equivalent to) $\lVert\mathbf{C}\rVert_{1,1}+\nu\sum_{i=1}^{I}\lVert\mathbf{U}_{i}\rVert_{\mathrm{F}}$. Unlike our model, the work [34] defines the regularizer as $\lVert\mathbf{C}\rVert_{1,1}+\nu\lVert\mathbf{U}\rVert_{1,r}$, where $r\in[1,+\infty)$, $\lVert\mathbf{U}\rVert_{1,r}=\sum_{i,j=1}^{d}\lVert\mathbf{U}[ij]\rVert_{r}$, and $\mathbf{U}[ij]=[\mathbf{U}_{1}[ij],...,\mathbf{U}_{I}[ij]]^{\top}\in\mathbb{R}^{I}$. The term $\lVert\mathbf{U}\rVert_{1,r}$ is a group lasso regularizer that can capture common sparsity structure among all local graphs. The second work [35] assumes that the common structure $\mathbf{C}$ is the average of all local graphs, i.e., $\mathbf{C}=\frac{1}{I}\sum_{i=1}^{I}\mathbf{A}_{i},\mathbf{U}_{i}=\mathbf{A}_{i}-\mathbf{C}$. A regularizer is then designed as $\lVert\mathbf{C}\rVert_{1,1}+\nu\sum_{i=1}^{I}\lVert\mathbf{U}_{i}\rVert_{1,1}$. Similar to our work, [36] proposes a regularizer $\lVert\mathbf{C}\rVert_{1,1}+\nu\sum_{i=1}^{I}\lVert\mathbf{U}_{i}\rVert_{\mathrm{F}}^{2}$. More recently, [37] designs two regularizers, i.e., $\sum_{i=1}^{I}\lVert\mathbf{U}_{i}\rVert_{1,1}$ and the other similar to [34].
Although closely related to these works, our model differs from them in the following aspects. (i) The regularizer $R(\mathbf{W})$ is different from existing works, which can implicitly assign a contribution weight for each local graph. This is used for the first time in the multiple graph topology inference problem. (ii) Our model and [36] learn multiple graphs based on the smoothness assumption (6), while [34, 35] are based on the Gaussian graphical model. (iii) We further consider the problem of learning graphs in the case of data silos, which has not been explored in the existing GL-related works. See Section IV for more details.
IV Model Optimization
In this section, we propose an algorithm with provable convergence analysis for solving (8) via the reformulation (13) in the target scenario.
IV-A The Proposed Algorithm
Our algorithm consists of two steps: (i) updating $\mathbf{w}_{i}$ in the local client $\mathcal{C}_{i},i=1,...,I$, and (ii) updating $\bm{\gamma}=[\gamma_{1},...,\gamma_{I}]^{\top}\in\mathbb{R}^{I}$ and $\mathbf{w}_{\mathrm{con}}$ in the central server. The complete flow is displayed in Algorithm 1.
Updating $\mathbf{w}_{i}$ in the local client $\mathcal{C}_{i}$: This update corresponds to lines 3-11 in Algorithm 1. In the $t$-th communication round, $\mathcal{C}_{i}$ first receives $\mathbf{w}^{(t)}_{\mathrm{con}}$ and $\gamma_{i}^{(t)}$ from the central server. Let $f_{i}^{(t)}(\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}):=g_{i}(\mathbf{w}_{i})+\frac{\rho\gamma_{i}^{(t)}}{2}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\lVert_{2}^{2}$, and the sub-problem of $\mathcal{C}_{i}$ becomes
$\mathbf{w}_{i}^{(t+1)}=$ $\underset{\mathbf{w}_{i}\in\mathcal{W}}{\mathrm{argmin}}\,\,f_{i}^{(t)}(\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}^{(t)})$
$=$ $\underset{\mathbf{w}_{i}\in\mathcal{W}}{\mathrm{argmin}}\,\frac{1}{N_{i}}\mathbf{z}_{i}^{\top}\mathbf{w}_{i}-\alpha\mathbf{1}^{\top}\log\left(\mathbf{S}\mathbf{w}_{i}+\zeta\mathbf{1}\right)$
$+2\beta\lVert\mathbf{w}_{i}\rVert_{2}^{2}+\frac{\rho\gamma_{i}^{(t)}}{2}\lVert\mathbf{w}_{i}-\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{2}^{2},$ (21)
where $\zeta$ is a small enough constant to avoid zero node degree. We use the accelerated projected gradient descent algorithm to update $\mathbf{w}_{i}$. At the beginning of local updates, $\mathcal{C}_{i}$ first initializes $\mathbf{w}_{i}^{(t,0)}=\mathbf{w}_{i}^{(t-1,K_{i}^{(t-1)})}$ and $\mathbf{w}_{i}^{(t,-1)}=\mathbf{w}_{i}^{(t-1,K_{i}^{(t-1)}-1)}$, where $K_{i}^{(t)}$ is the number of local loops of $\mathcal{C}_{i}$ in the $t$-th outer iteration. Besides, $\mathbf{w}_{i}^{(t,k)}$ represents the updated graph of $\mathcal{C}_{i}$ in the $t$-th outer iteration and the $k$-th local loop. We then update the $i$-th local graph $K_{i}^{(t)}$ times by
$\mathbf{w}_{i,\mathrm{ex}}^{(t,k)}=\mathbf{w}_{i}^{(t,k)}+\xi\left(\mathbf{w}_{i}^{(t,k)}-\mathbf{w}_{i}^{(t,k-1)}\right)$ (22)
$\breve{\mathbf{w}}_{i}^{(t,k+1)}=\mathbf{w}_{i,\mathrm{ex}}^{(t,k)}-\eta_{w}\bigg{(}\nabla g_{i}(\mathbf{w}_{i,\mathrm{ex}}^{(t,k)})+\rho\gamma_{i}^{(t)}\left(\mathbf{w}_{i,\mathrm{ex}}^{(t,k)}-\mathbf{w}_{\mathrm{con}}^{(t)}\right)\bigg{)}$ (23)
$\mathbf{w}_{i}^{(t,k+1)}=\mathrm{Proj}_{\mathcal{W}}\left(\breve{\mathbf{w}}_{i}^{(t,k+1)}\right),$ (24)
where $\xi\in[0,1)$ is an momentum weight. In (23), $\nabla g_{i}(\mathbf{w})$ is calculated as $\frac{1}{N_{i}}\mathbf{z}_{i}-\alpha\mathbf{S}^{\top}\left(\frac{1}{\mathbf{S}\mathbf{w}+\zeta\mathbf{1}}\right)+4\beta\mathbf{w}$. Moreover, $\eta_{w}$ is the stepsize, the choice of which will be discussed in the next subsection. The operator $\mathrm{Proj}_{\mathcal{W}}\left(\cdot\right)$ means projecting variables into ${\mathcal{W}}$. When local updates finish, $\mathcal{C}_{i}$ sends $\mathbf{w}_{i}^{(t+1)}=\mathbf{w}_{i}^{(t,K_{i}^{(t)})}$ to the central server. Note that the local update of $\mathbf{w}_{i}$ is inexact in one communication round since we run (22)-(24) for $K_{i}^{(t)}$ times without convergence. We use inexact updates because we aim to update all local graphs synchronously. Due to system heterogeneity, the time required for each client to update $\mathbf{w}_{i}$ until convergence may vary widely. It takes longer if the central server waits for all clients to update their graphs until convergence.
Updating $\mathbf{w}_{\mathrm{con}}$ and $\bm{\gamma}$ in the central server: The updates correspond to lines 12-14 in Algorithm 1. After the central server receives all the updated local graphs, $\mathbf{w}_{1}^{(t+1)},...,\mathbf{w}_{I}^{(t+1)}$, we update $\mathbf{w}_{\mathrm{con}}^{(t+1)}$ by solving the following problem
$\mathbf{w}_{\mathrm{con}}^{(t+1)}$ $=\underset{\mathbf{w}_{\mathrm{con}}\in\mathcal{W}}{\mathrm{argmin}}\,\,\sum_{i=1}^{I}f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}_{\mathrm{con}})+\lambda\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}$
$=\underset{\mathbf{w}_{\mathrm{con}}\in\mathcal{W}}{\mathrm{argmin}}\,\sum_{i=1}^{I}\frac{\rho\gamma_{i}^{(t)}}{2}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{\mathrm{con}}\rVert_{2}^{2}+\lambda\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}.$ (25)
The problem is equivalent to
$\mathbf{w}_{\mathrm{con}}^{(t+1)}=$ $\underset{\mathbf{w}_{\mathrm{con}}\in\mathcal{W}}{\mathrm{argmin}}\,\frac{1}{2}\left\lVert\mathbf{w}_{\mathrm{con}}-\frac{\sum_{i=1}^{I}\gamma_{i}^{(t)}\mathbf{w}_{i}^{(t+1)}}{\sum_{i=1}^{I}\gamma_{i}^{(t)}}\right\rVert_{2}^{2}$
$+\frac{\lambda}{\rho\sum_{i=1}^{I}\gamma_{i}^{(t)}}\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}.$ (26)
By defining $C_{\gamma}^{(t)}:=\sum_{i=1}^{I}\gamma_{i}^{(t)}$ and $\mu^{(t)}:=\frac{\lambda}{C_{\gamma}^{(t)}\rho}$, we obtain
$\mathbf{w}_{\mathrm{con}}^{(t+1)}=\mathrm{prox}_{\mu^{(t)}\lVert\cdot\rVert_{1}}\left(\frac{\sum_{i=1}^{I}\gamma_{i}^{(t)}\mathbf{w}_{i}^{(t+1)}}{C_{\gamma}^{(t)}}\right),$ (27)
where $\mathrm{prox}_{\mu^{(t)}\lVert\cdot\rVert_{1}}(\cdot)$ is the proximal operator of $\ell_{1}$ norm. It is observed from (27) that $\mathbf{w}_{\mathrm{con}}^{(t+1)}\in\mathcal{W}$ since it is a combination of $\mathbf{w}_{i}^{(t+1)}$ with positive weights, and the proximal operator will not move the vector out of $\mathcal{W}$.
After obtaining $\mathbf{w}_{\mathrm{con}}^{(t+1)}$, we update $\gamma_{i}^{(t+1)}$ as
$\gamma_{i}^{(t+1)}=\frac{1}{\left\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{\mathrm{con}}^{(t+1)}\right\rVert_{2}+\epsilon_{\gamma}},\,\,\text{for}\,\,i=1,...,I,$ (28)
where $\epsilon_{\gamma}$ is a small enough constant. Finally, we send $\mathbf{w}_{\mathrm{con}}^{(t+1)}$ and $\gamma_{i}^{(t+1)}$ back to the client $\mathcal{C}_{i}$.
Algorithm 1 The algorithm for solving (8)
0:  $\alpha$, $\beta$, $\nu$, $\xi$, $\lambda$, and signals $\mathbf{X}_{1}$,…, $\mathbf{X}_{I}$
1:  Initialize $\gamma_{i}^{(0)}=1/I$, $\mathbf{w}_{\mathrm{con}}^{(0)}=\mathbf{w}_{i}^{(0)}=\mathbf{w}_{i}^{(-1,0)}=\mathbf{w}_{i}^{(-1,-1)}$ for $i=1,...,I$, and let $K_{i}^{(-1)}=0$
2:  for $t=0,...,T-1$ do
3:     / /  Update $\mathbf{w}_{1},...,\mathbf{w}_{I}$ in parallel in local clients
4:     for $i=1,...,I$ in parallel do
5:        Receive $\gamma_{i}^{(t)}$ and $\mathbf{w}^{(t)}_{\mathrm{con}}$ from the central server
6:        Initialize $\mathbf{w}_{i}^{(t,0)}=\mathbf{w}_{i}^{(t-1,K_{i}^{(t-1)})}$ and $\mathbf{w}_{i}^{(t,-1)}=\mathbf{w}_{i}^{(t-1,K_{i}^{(t-1)}-1)}$
7:        for $k=0,...,K_{i}^{(t)}-1$ do
8:           Update $\mathbf{w}_{i}^{(t,k+1)}$ using (22)-(24)
9:        end for
10:        Let $\mathbf{w}_{i}^{(t+1)}=\mathbf{w}_{i}^{(t,K_{i}^{(t)})}$ and send it to central server
11:     end for
12:     / /   Update $\mathbf{w}_{\mathrm{con}}$ and $\bm{\gamma}$ in central server
13:     Update $\mathbf{w}_{\mathrm{con}}^{(t+1)}$ using (27)
14:     Update ${\gamma}_{i}^{(t+1)}$ using (28), for $i=1,...,I$
15:     Send $\mathbf{w}_{\mathrm{con}}^{(t+1)}$ and $\gamma_{i}^{(t+1)}$ to the $i$-th client $\mathcal{C}_{i}$
16:  end for
17:  return  $\mathbf{w}_{1}^{(T)}$, …,$\mathbf{w}_{I}^{(T)}$ and $\mathbf{w}_{\mathrm{con}}^{(T)}$
IV-B Convergence Analysis
Before proceeding with the convergence analysis, let us make some technical assumptions.
Assumption 3.
The feasible set is convex and compact, i.e., the updated graphs in the search space are bounded. For any updated $\mathbf{y}$ and $\mathbf{y}^{\prime}\in{\mathcal{W}}$, there exists a constant such that $\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert_{2}\leq B_{r}$. Furthermore, for any updated $\mathbf{y}$ and $\mathbf{y}^{\prime}\in\widetilde{\mathcal{W}}$ , $\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert_{2}\leq B_{w}$, where $\widetilde{\mathcal{W}}=\mathrm{conv}[\cup_{0\leq\xi\leq 1}\{\mathbf{y}+\xi(\mathbf{y}-\mathbf{y}^{\prime})||\;\mathbf{y},\mathbf{y}^{\prime}\in\mathcal{W}\}]$ is the feasible set extended by the momentum update.
Assumption 3 is common in constrained optimization problems [38, 39]. It implies that the weights $\gamma_{i}^{(t)}$ updated via (28) are bounded in $\left[\frac{1}{2B_{r}+\epsilon_{\gamma}},\frac{1}{\epsilon_{\gamma}}\right]$. The boundary is usually not reached, and we define two positive constants $\gamma_{\min}$ and $\gamma_{\max}$ to represent the possible minimum and maximum values of $\gamma_{i}^{(t)}$ for all $i$ and $t$. Next, we have the following proposition.
Proposition 1.
The objective function $f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}_{\mathrm{con}})$ is $\rho\gamma_{i}^{(t)}$-Lipschitz smooth w.r.t. $\mathbf{w}_{\mathrm{con}}$, and the function $f_{i}^{(t)}(\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}^{(t)})$ is $L_{i}^{(t)}$-Lipschitz smooth w.r.t. $\mathbf{w}_{i}$ on $\widetilde{\mathcal{W}}$, where $L_{i}^{(t)}:=4\beta+{2\alpha(d-1)}/{\zeta^{2}}+\rho\gamma_{i}^{(t)}$.
Proof:
We place the proof in Appendix B. ∎
We further make the following assumption to determine the stepsize in Algorithm 1.
Assumption 4.
The (constant) stepsize satisfies $\eta_{w}\leq 1/L_{max}$, where $L_{max}$ is the maximum value of $L_{i}^{(t)}$ for all $i$ and $t$, i.e., $L_{max}=4\beta+{2\alpha(d-1)}/{\zeta^{2}}+\rho\gamma_{\max}$.
This assumption is necessary for gradient descent-type algorithms. Finally, we provide the convergence analysis of our proposed algorithm as follows.
Theorem 2.
Suppose that Assumptions 3 and 4 hold, the sequences $\mathbf{w}_{i}^{(t)}$, $\mathbf{w}_{\mathrm{con}}^{(t)}$ generated from our algorithm satisfy
$\frac{1}{T}\sum_{t=0}^{T-1}\left(\sum_{i=1}^{I}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}\right)\leq\frac{C_{1}}{T},$ (29)
$\frac{1}{T}\sum_{t=0}^{T-1}\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{2}^{2}\leq\frac{C_{2}}{T},$ (30)
where
$C_{1}=\frac{2\Upsilon\eta_{w}}{1-\xi^{2}},\;\;\;C_{2}=\frac{2\Upsilon}{\rho I\gamma_{\min}},$
$\Upsilon=\sum_{i=1}^{I}f_{i}^{(0)}\left(\mathbf{w}_{i}^{(0)},\mathbf{w}^{(0)}_{\mathrm{con}}\right)-f_{i}^{*}+\lambda\sqrt{p}B_{r}.$ (31)
In (31), $f_{i}^{*}$ denotes the minimum value of $f_{i}(\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}})$.
Proof:
See Appendix C for details. ∎
The theorem reveals that as $T$ increases, the average cumulative error between two consecutive iterations will decay to zero for both $\mathbf{w}_{i}$ and $\mathbf{w}_{\mathrm{con}}$, indicating that our algorithm enjoys a convergence rate of $\mathcal{O}(1/T)$.
IV-C Privacy Analysis
In our algorithm, each client $\mathcal{C}_{i}$ uses its private data $\mathbf{X}_{i}$ to update the local graph $\mathbf{w}_{i}$ by solving (21). After local updates, clients send model updates instead of raw data to the central server to update the consensus graph and weights. Following the paradigm of FL, all data is kept locally without any leakage during execution, which protects data privacy to a certain extent. Therefore, the proposed algorithm can learn graphs in the target scenario. However, inference attacks [40] and model inversion attacks [41] show that the updates sent by local clients may still reveal private information. To further prevent information leakage, some popular techniques such as differential privacy [42] may be helpful. This is beyond the scope of this study and will be left to future work.
IV-D Some Discussions
Our algorithm differs from the existing (personalized) FL algorithms in the following ways due to the specificity of the GL problem. (i) Unlike the algorithms of cross-device FL [25, 14], our algorithm learns graphs in a cross-silo FL scenario. Specifically, all clients, rather than a subset of clients, participate in local updates in each communication round. In addition, local clients utilize all stored data to update graphs instead of sampling a batch of data like many cross-device FL algorithms [25, 14]. (ii) The optimization problems corresponding to local and global updates are customized. For example, our algorithm updates the consensus graph by solving a $\ell_{1}$ norm regularized quadratic optimization problem, whereas the global update of existing algorithms is a simple aggregation of local models [25, 17, 27]. (iii) Finally, compared to decentralized PFL algorithms [26, 28], our algorithm follows the “central server-clients” schema.
V Experiments
In this section, we will test our proposed framework using both synthetic data and real-world data. First, some experimental setups are introduced.
V-A Experimental Setups
TABLE I: Performance of varying data sizes.
  $N=20$ $N=50$ $N=100$
$\mathrm{Precision}\uparrow$ $\mathrm{Recall}\uparrow$ $\mathrm{FS}\uparrow$ $\mathrm{RE}\downarrow$ $\mathrm{Precision}\uparrow$ $\mathrm{Recall}\uparrow$ $\mathrm{FS}\uparrow$ $\mathrm{RE}\downarrow$ $\mathrm{Precision}\uparrow$ $\mathrm{Recall}\uparrow$ $\mathrm{FS}\uparrow$ $\mathrm{RE}\downarrow$
IGL 0.594 0.469 0.521 0.915 0.692 0.620 0.651 0.778 0.776 0.739 0.755 0.698
FedAvg 0.434 0.759 0.552 0.802 0.464 0.792 0.584 0.768 0.519 0.789 0.624 0.744
FedProx 0.423 0.767 0.544 0.765 0.523 0.728 0.606 0.718 0.623 0.703 0.655 0.690
MRMTL 0.522 0.551 0.547 0.881 0.647 0.688 0.663 0.756 0.738 0.789 0.760 0.680
Ditto 0.582 0.547 0.557 0.884 0.662 0.671 0.663 0.775 0.732 0.780 0.750 0.696
PPGL-L 0.547 0.579 0.561 0.911 0.625 0.760 0.678 0.731 0.698 0.870 0.774 0.619
PPGL-C 0.636 0.516 0.566 0.874 0.781 0.656 0.710 0.701 0.881 0.736 0.800 0.627
  • 1
    In this paper, $\uparrow$ indicates that larger values are better, and $\downarrow$ indicates that smaller values are better.
TABLE II: Performance of varying data heterogeneity.
  $q=0.3$ $q=0.6$ $q=0.9$
$\mathrm{Precision}\uparrow$ $\mathrm{Recall}\uparrow$ $\mathrm{FS}\uparrow$ $\mathrm{RE}\downarrow$ $\mathrm{Precision}\uparrow$ $\mathrm{Recall}\uparrow$ $\mathrm{FS}\uparrow$ $\mathrm{RE}\downarrow$ $\mathrm{Precision}\uparrow$ $\mathrm{Recall}\uparrow$ $\mathrm{FS}\uparrow$ $\mathrm{RE}\downarrow$
IGL 0.701 0.593 0.640 0.783 0.736 0.645 0.685 0.758 0.822 0.706 0.759 0.717
FedAvg 0.385 0.841 0.527 0.835 0.460 0.867 0.599 0.783 0.811 0.922 0.862 0.672
FedProx 0.437 0.720 0.543 0.770 0.569 0.750 0.645 0.716 0.874 0.863 0.868 0.641
MRMTL 0.662 0.691 0.674 0.761 0.690 0.730 0.707 0.753 0.816 0.794 0.807 0.663
Ditto 0.667 0.675 0.668 0.765 0.698 0.714 0.704 0.739 0.821 0.783 0.804 0.667
PPGL-L 0.637 0.736 0.679 0.740 0.661 0.776 0.715 0.700 0.800 0.812 0.805 0.635
PPGL-C 0.462 0.516 0.487 0.874 0.653 0.726 0.687 0.752 0.825 0.927 0.873 0.634
1) Graph generation: We first generate Gaussian radial basis function (RBF) graphs $\mathcal{G}_{0}$ by following the method in [6]. Specifically, we generate 20 vertices whose coordinates are randomly in a unit square. Edge weights are calculated using $\mathrm{exp}(-\mathrm{dist}(i,j)^{2}/2\sigma^{2}_{r})$, where $\mathrm{dist}(i,j)$ is the distance between vertices $i$ and $j$, and $\sigma_{r}=0.5$ is a kernel width parameter. We remove the edges whose weights are smaller than $0.7$. Then, we keep a certain proportion—say $q$—of edges in $\mathcal{G}_{0}$ unchanged to form the consensus graph. Finally, we add $(1-q)|\mathcal{E}_{0}|$ edges randomly to the consensus graph to form heterogeneous graphs $\mathcal{G}_{1},....,\mathcal{G}_{I}$, where $|\mathcal{E}_{0}|$ is the number of edges in $\mathcal{G}_{0}$. We denote $1-q$ as the degree of heterogeneity.
2) Signal generation: For the $i$-th local graph $\mathcal{G}_{i}$, we generate $N_{i}$ smooth signals from the distribution [6]
$\mathbf{X}_{i,n}\sim\mathcal{N}\left(\mathbf{0},(\mathbf{L}_{i}^{*})^{{\dagger}}+\sigma_{w}^{2}\mathbf{I}\right),\;\;n=1,...,N_{i},$ (32)
where $\sigma_{w}$ is noise scale and $\mathbf{L}_{i}^{*}$ is the real Laplacian matrix of $\mathcal{G}_{i}$, and $\mathbf{X}_{i,n}$ is the $n$-th data (column) of $\mathbf{X}_{i}$.
3) Evaluation metric: Four metrics are employed to evaluate the learned graphs, as listed in (33). The first three metrics of classification tasks since determining whether two vertices are connected can be regarded as a binary classification problem. Specifically, $\mathrm{TP}$ is the true positive rate, $\mathrm{TN}$ is the true negative rate, $\mathrm{FP}$ is the false positive rate and $\mathrm{FN}$ denotes false negative rate. The third metric, F1-score ($\mathrm{FS}$) is the harmonic mean of $\mathrm{Precision}$ and $\mathrm{Recall}$. The last metric is the relative error ($\mathrm{RE}$), where $\widehat{\mathbf{w}}$ is the learned graph, and $\mathbf{w}^{*}$ is the groundtruth. The results of all metrics for local graphs are the average of $I$ clients.
$\mathrm{Precision}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},\;\;\;\;\mathrm{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$
$\mathrm{FS}=\frac{2\mathrm{TP}}{2\mathrm{TP}+\mathrm{FN}+\mathrm{FP}},\;\;\;\mathrm{RE}=\frac{\lVert\widehat{\mathbf{w}}-\mathbf{w}^{*}\rVert_{2}}{\lVert\mathbf{w}^{*}\rVert_{2}}.$ (33)
4) Baselines: We employ three categories of baselines. (i) We learn all local graphs independently (IGL), which enjoys full privacy as local clients do not release any data, including model updates. (ii) We use two traditional FL algorithm, FedAvg [14] and FedProx [43], to solve the following problem
$\underset{\mathbf{w}\in\mathcal{W}}{\min}$ $\frac{1}{N^{\prime}}\sum_{i=1}^{I}{\mathbf{z}_{i}^{\top}\mathbf{w}-\alpha\mathbf{1}^{\top}\log\left(\mathbf{S}\mathbf{w}+\zeta\mathbf{1}\right)+2\beta\lVert\mathbf{w}\rVert_{2}^{2}},$ (34)
where $N^{\prime}=\sum_{i=1}^{I}N_{i}$. The model learns a global graph collaboratively using data from all clients regardless of data heterogeneity. (iii) We leverage two personalized FL algorithms, MRMTL [44] and Ditto [45], to jointly learn all local graphs by further considering data heterogeneity. For our algorithm, we use PPGL-L and PPGL-C to represent the learned local and consensus graphs, respectively.
5) Determination of parameters: For our model, $\alpha$ is fixed as $1$, and $\beta$ is selected as the value achieving the best $\mathrm{FS}$ in IGL. Parameters $\nu$ and $\lambda$ are determined by grid search in $[1,100]$ and $[0.01,1]$. On algorithmic side, we set $\xi=0.9,K_{i}^{(t)}=K=5,T=50$. The stepsize $\eta_{w}$ is $0.005$, which is small enough to satisfy $\eta_{w}\leq 1/L_{max}$. Moreover, we let $I=5$ and $\sigma_{w}=0.1$ for synthetic data. All parameters of baselines are selected as those achieving the best $\mathrm{FS}$ values.
V-B Synthetic Data
1) Data size: We first test the effect of data size. We fix $q$ as 0.5 and assume that all clients have the same data size, i.e., $N_{i}=N,i=1,...,I$. We vary $N$ from 20 to 100, and the results are shown in Table.I. It is observed that FedAvg and FedProx tend to obtain high $\mathrm{Recall}$ but low $\mathrm{Precision}$, meaning that the corresponding graphs contain more edges. As $N$ increases, FedAvg yields the worst $\mathrm{FS}$, which is a comprehensive metric for evaluating the learned graph topology. The reason is that FedAvg learns a global graph using data from all clients. However, local graphs are heterogeneous in our experiment. Therefore, the learned global graph is far from local graphs. In sharp contrast, IGL learns local graphs independently and achieves better performance than FedAvg when $N$ is large. Our model also learns a personalized graph for each local client. Unlike IGL, we learn all local graphs jointly via a consensus graph, leading to better performance than IGL. Our model is also superior to Ditto and MRMTL—two PFL algorithms—since it is tailored for graph learning problems. Additionally, our model learns a consensus graph that captures the common structures of local graphs.
We also test our model with different data sizes for different clients. We set $N_{i}$ to 20, 40, 60, 80, and 100 for $i=1,...,5$, respectively. As illustrated in Fig.2, clients with larger $N_{i}$ perform better than those with small $N_{i}$. However, our model can improve the performance of clients with small $N_{i}$ because they can “borrow” information from the clients with large $N_{i}$.
2) Data heterogeneity: We next explore the effect of data heterogeneity. In this experiment, we fix $N_{i}=50$ for all clients and vary $q$ from 0.3 to 0.9. As listed in Table II, when data heterogeneity is small (large $q$), FedAvg and FedProx achieve satisfactory performance since the consensus graph is close to local graphs. In this case, FedAvg and FedProx benefit from using all data to learn a global graph. However, with the increase in data heterogeneity, the personalized algorithms outperform FedAvg and FedProx. Our model obtains the best performance for all data heterogeneity since we learn personalized graphs and a consensus graph simultaneously.
Figure 2: The results of varying data sizes of different clients. The data sizes $N_{1},...,N_{5}$ are 20, 40, 60, 80, and 100, respectively.
3) Adaptive weights: Then, we test the effectiveness of the method of adjusting weights. Two factors are considered, i.e., data size and data heterogeneity. For data size, we let $q$ be 0.5 and set $N_{i}$ to 20, 40, 60, 80, and 100 for $\mathcal{C}_{1}-\mathcal{C}_{5}$, respectively. For data heterogeneity, we fix $N_{i}=100$ and first generate a graph $\mathcal{G}_{1}$ for $\mathcal{C}_{1}$. The graphs of $\mathcal{C}_{2}$-$\mathcal{C}_{5}$ are generated based on $\mathcal{G}_{1}$ with $q$ equal to 0.8, 0.6, 0.4, and 0.2, respectively. The results are displayed in Table III, and two trends can be observed. First, local graphs with large data sizes contribute more to the consensus graph. Second, local graphs close to the consensus graph obtain larger weights. Therefore, the weight adjustment strategy can effectively allocate a reasonable weight for each client.
TABLE III: The learned $\bm{\gamma}$ of different clients
Index of clients 1 2 3 4 5
Data size 0.541 0.584 0.661 0.707 0.718
Graph heterogeneity 0.837 0.808 0.720 0.735 0.584
4) Parameter sensitivity: In this experiment, we set $N_{i}=100$ and $q=0.5$, and evaluate the learning performance of different $\lambda$ and $\nu$. As displayed in Fig.3, local graphs are less sensitive to the parameters than the consensus graph. When $\nu$ and $\lambda$ are too large, the performance of the consensus graph degrades rapidly. However, there exist combinations of $\nu$ and $\lambda$ that achieve the highest $\mathrm{FS}$ for both local graphs and the consensus graph. We need to grid search for the optimal combination of $\nu$ and $\lambda$, especially for consensus graphs.
(a) Local graphs
(b) The consensus graph
Figure 3: The results of parameter sensitivity.
5) Convergence: Finally, we test the convergence of the proposed algorithms. Fig.4 shows that our algorithms with different $K$ can converge as stated in Theorem 2. When $K$ is small, fewer local updates are performed. In this case, more communication rounds between local clients and the central server are required. On the other hand, when $K$ is large, our algorithm only takes a few rounds to converge.
(a) Local graphs
(b) The consensus graph
Figure 4: The convergence of the proposed algorithm.
V-C Real-world Data
1) COIL-20 data: We employ the COIL-20 dataset222 https://www.cs.columbia.edu/CAVE/software/softlib/coil-20.php, which is a collection of gray-scale images including 20 objects taken from 360 degrees, to learn the relationships between these images. Each object has 72 images (five degrees an image) of size $32\times 32$. We randomly select four objects and divide the images of each object into three views. Each view contains images taken in consecutive 60 degrees, e.g., $[0^{\circ},55^{\circ}]$. Therefore, each view contains 48 images taken from 4 objects, i.e., 12 images per object. A basic assumption is that the 48 images of different views are defined on the same node set. The image itself is taken as graph signals, i.e., $\mathbf{X}_{i}\in\mathbb{R}^{48\times 1024}$. Data distributions of different views are heterogeneous because they are from different shooting angles. Furthermore, we assume that photos from three views are taken by three photographers, and they are reluctant to share their photos, which is the concern of this study. Our goal is to learn a graph representing the relationships among these 48 images for each view, i.e., $I=3$, under data silos. The three graphs should share some common structures because they come from the same objects. Moreover, there should be four communities in the relation graphs since all images belong to four objects. Therefore, to evaluate the learned graphs, we employ the Louvain algorithm [46] to detect communities in the learned graphs. Three metrics are leveraged to evaluate the detection results, i.e., normalized mutual information ($\mathrm{NMI}$), Fowlkes and Mallows index ($\mathrm{FMI}$) [47], and Rand Index ($\mathrm{RI}$). The labels of images are taken as the ground truth, and the results of local graphs are the average of three views. As displayed in Table IV, the Louvain algorithm successfully finds all communities in the consensus graph learned by our models. Instead, some mistakes are made in the graphs learned by IGL and FedAvg. This is expected since the consensus graph captures the common structure of local graphs, which may remove noisy edges of different views. Furthermore, the local graphs learned by our model also outperform those of other models since our proposed model is tailored for graph learning problems. As shown in Fig.5, the communities can be clearly observed in the consensus graph. However, more confusing edges appear in the graphs of IGL and FedAvg.
TABLE IV: The community detection results of graphs learned by different methods.
IGL FedAvg FedProx MRMTL Ditto PPGL-L PPGL-C
$\mathrm{NMI}\uparrow$ 0.779 0.745 0.767 0.795 0.786 0.865 1
$\mathrm{RI}\uparrow$ 0.855 0.843 0.851 0.891 0.883 0.919 1
$\mathrm{FMI}\uparrow$ 0.755 0.736 0.745 0.773 0.768 0.852 1
(a) IGL- $\mathcal{C}_{2}$
(b) FedAvg
(c) PPGL-L- $\mathcal{C}_{2}$
(d) PPGL-C
Figure 5: The figure depicts the communities in the learned graphs. The suffix $-\mathcal{C}_{2}$ stands for the local graphs of $\mathcal{C}_{2}$.
(a) Sagittal
(b) Axial
(c) Coronal
(d) Sagittal
(e) Axial
(f) Coronal
Figure 6: The figure depicts our learned graphs from three views. The top and bottom rows show the graphs of non-autistic and autistic subjects, respectively. The red edges are those in the consensus graph, while the blue edges are those unique edges in local graphs.
2) Medical data: We finally employ blood-oxygenation-level-dependent (BOLD) time series extracted from fMRI data to learn brain functional connectivity graphs. The basic assumption is that autism may affect brain functional connectivity. If we can learn about the differences in brain functional connectivity graphs between autistic and non-autistic people, it may help us better understand the mechanism of autism. However, medical data are privacy-sensitive, and their transmission to an unreliable central server is usually prohibited. In this experiment, the used dataset [^3] http://preprocessed-connectomes-project.org/abide/ contains 539 autistic individuals and 573 typical controls, from which we select three autistic and three non-autistic subjects. Each client (subject) contains 176 signals (the length of the BOLD time series is 176). Besides, we select 34 functional regions of interest from 90 standard regions of the Anatomical Automatic Labeling (AAL) template, indicating that $\mathbf{X}_{i}\in\mathbb{R}^{34\times 176},i=1,...,6$. Figure 6 depicts the learned graphs using our model. It is observed that graphs of autistic and non-autistic individuals share many common edges that are learned in the consensus graph. On the other hand, there are some noticeable topological changes between the graphs of autistic and non-autistic individuals. These edges reflect changes in the functional connectivity caused by autism, which may aid in diagnosing autism. Furthermore, the graphs of autistic individuals display fewer edges, which is consistent with the current study [48]. In the future, a domain expert may help interpret the results from a medical perspective.
VI Conclusion
In this paper, we proposed a framework to learn graphs under data silos. In our framework, we jointly learned a personalized graph for each client to handle data heterogeneity, and a consensus graph for all clients to capture global information. Then, we devised a tailored algorithm in which all private data are processed in local clients to preserve data privacy. Convergence analyses for the proposed algorithm were provided. Extensive experiments showed that our approach can efficiently learn graphs in the target scenario. Future research may include generalizing our framework to other graph learning models except for the smoothness assumption.
Appendix A Proof of Theorem 1
For $\mathbf{\Delta}\in\mathbb{R}^{p\times(I+1)}$, we define $\phi(\mathbf{\Delta})=G(\mathbf{W}^{*}+\mathbf{\Delta})-G(\mathbf{W}^{*})+\lambda({R}(\mathbf{W}^{*}+\mathbf{\Delta})-{R}(\mathbf{W}^{*}))$. Based on the optimality of the solution, we have $\phi(\widehat{\mathbf{\Delta}})\leq 0$, where $\widehat{\mathbf{\Delta}}=\widehat{\mathbf{\mathbf{W}}}-\mathbf{\mathbf{W}}^{*}$. Denote the support space of $\mathbf{W}^{*}$ as :=$\mathcal{M}:=\{(i,j):\mathbf{W}^{*}[ij]\neq 0,i=1,...p,j=1,...,I+1\}$. The orthogonal complement of $\mathcal{M}$ is defined as the set :=$\mathcal{M}^{\perp}:=\{\mathbf{W}^{\prime}\in\mathbb{R}^{p\times{(I+1)}}:\left\langle\mathbf{W},\mathbf{W}^{\prime}\right\rangle=0,\forall\mathbf{W}\in\mathcal{M}\}$. For any $\mathbf{Y}\in\mathbb{R}^{p\times{(I+1)}}$, we have $\mathbf{Y}=\mathbf{Y}_{\mathcal{M}}+\mathbf{Y}_{\mathcal{M}^{\perp}}$, where $\mathbf{Y}_{\mathcal{M}}$ and $\mathbf{Y}_{\mathcal{M}^{\perp}}$ are the projection of $\mathbf{Y}$ into $\mathcal{M}$ and $\mathcal{M}^{\perp}$, respectively. Then, we provide some Lemmas essential to our proof.
Lemma 1.
Properties of $R(\mathbf{W})$
(i) The regularizer $R(\mathbf{W})$ is convex and decomposable w.r.t. $(\mathcal{M},\mathcal{M}^{\perp})$, i.e.,
$R(\mathbf{W}_{1}+\mathbf{W}_{2})=R(\mathbf{W}_{1})+R(\mathbf{W}_{2}),\;\forall\mathbf{W}_{1}\in\mathcal{M},\mathbf{W}_{2}\in\mathcal{M}^{\perp}.$ (35)
Moreover,
$R(\mathbf{W}^{*}+\mathbf{\Delta})-R(\mathbf{W}^{*})\geq R(\mathbf{\Delta}_{\mathcal{M}^{\perp}})-R(\mathbf{\Delta}_{\mathcal{M}}).$ (36)
(ii) For $\mathbf{W}\in\mathcal{M}$, $R(\mathbf{W})$ satisfies:
$R(\mathbf{W})\leq(\sqrt{s}+\nu\sqrt{Ip\omega_{\max}(\mathbf{L}_{m})})\lVert\mathbf{W}\rVert_{\mathrm{F}}.$ (37)
Proof:
Proof of Lemma 1
(i) The convexity of $R(\mathbf{W})$ comes from the convexity of $\ell_{1}$ and $\ell_{2}$ norm. The linear transformation of the columns of $\mathbf{W}$ does not change convexity. According to the definition of decomposition (35), it is not difficult to check that $R(\mathbf{W})$ is decomposable. Moreover, let $\mathbf{W}^{*}+\mathbf{\Delta}=\mathbf{W}^{*}+\mathbf{\Delta}_{\mathcal{M}^{\perp}}+\mathbf{\Delta}_{\mathcal{M}}$. We have $R(\mathbf{W}^{*}+\mathbf{\Delta})\geq R(\mathbf{W}^{*}+\mathbf{\Delta}_{\mathcal{M}^{\perp}})-R(\mathbf{\Delta}_{\mathcal{M}})$ by triangle inequality. Finally, we yield
$R(\mathbf{W}^{*}+\mathbf{\Delta})-R(\mathbf{W}^{*})$
$\geq$ $R(\mathbf{W}^{*}+\mathbf{\Delta}_{\mathcal{M}^{\perp}})-R(\mathbf{\Delta}_{\mathcal{M}})-R(\mathbf{W}^{*})$
$=$ $R(\mathbf{\Delta}_{\mathcal{M}^{\perp}})-R(\mathbf{\Delta}_{\mathcal{M}}).$ (38)
(ii) For $\mathbf{W}\in\mathcal{M}$, we have
$\frac{R(\mathbf{W})}{\lVert\mathbf{W}\rVert_{\mathrm{F}}}\leq\underset{\mathbf{W}\in\mathcal{M}}{\sup}\frac{\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}}{\lVert\mathbf{W}\rVert_{\mathrm{F}}}+\underset{\mathbf{W}\in\mathcal{M}}{\sup}\frac{\nu\sum_{i=1}^{I}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}}{\lVert\mathbf{W}\rVert_{\mathrm{F}}}$
$\leq$ $\underset{\mathbf{W}\in\mathcal{M}}{\sup}\frac{\lVert\mathbf{W}\rVert_{1,1}}{\lVert\mathbf{W}\rVert_{\mathrm{F}}}+\underset{\mathbf{W}\in\mathcal{M}}{\sup}\frac{\nu\sqrt{Ip}\sqrt{\sum_{i=1}^{I}\lVert\mathbf{w}_{i}-\mathbf{w}_{\mathrm{con}}\rVert_{2}^{2}}}{\lVert\mathbf{W}\rVert_{\mathrm{F}}}$
$=$ $\sqrt{s}+\underset{\mathbf{W}\in\mathcal{M}}{\sup}\frac{\nu\sqrt{Ip}\sqrt{\sum_{j=1}^{p}\widetilde{\mathbf{w}}_{j}\mathbf{L}_{m}\widetilde{\mathbf{w}}_{j}^{\top}}}{\lVert\mathbf{W}\rVert_{\mathrm{F}}}$
$\leq$ $\sqrt{s}+\frac{\nu\sqrt{Ip}\sqrt{\sum_{j=1}^{p}\lVert\mathbf{L}_{m}\rVert_{2}\lVert\widetilde{\mathbf{w}}_{j}\rVert_{2}^{2}}}{\lVert\mathbf{W}\rVert_{\mathrm{F}}}$
$=$ $\sqrt{s}+\nu\sqrt{Ip\omega_{\max}(\mathbf{L}_{m})},$ (39)
where $\widetilde{\mathbf{w}}_{j}=\left[\mathbf{W}[j1],...,\mathbf{W}[jI]\right]\in\mathbb{R}^{I+1}$ is the $j$-th row vector of $\mathbf{W}$. Finally, we complete the proof of Lemma 1.
Lemma 2.
Properties of $G(\mathbf{W})$
(i) Suppose $R^{*}(\cdot)$ is the dual norm of $R(\cdot)$, then we have $R^{*}(\nabla G(\mathbf{W}^{*}))\leq\kappa$ with probability at least $1-\frac{4I}{d}$, where
$\kappa=$ $*\displaystyle\omega_{\min}(\mathbf{L}_{m})\sqrt{I+1}\Bigg{(}\sqrt{\frac{2d(d-1)}{B_{d}}}+4\beta B_{w^{*}}$
$+160\sqrt{2}\max_{i,k}\varpi(\mathbf{L}_{i}^{*}){[kk]}\sqrt{\frac{\ln d}{N_{\min}}}+4\max_{i}\left\lVert\varpi(\mathbf{L}_{i}^{*})\right\rVert_{\infty}\Bigg{)}.$ (40)
(ii) The function $G(\mathbf{W})$ satisfies
$G(\mathbf{W}^{*}+\mathbf{\Delta})-G(\mathbf{W}^{*})-\left\langle\nabla G(\mathbf{W}^{*}),\mathbf{\Delta}\right\rangle\geq 2\beta\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}^{2}.$ (41)
Proof:
Proof of Lemma 2.
(i)Taking the derivative of $G(\mathbf{W})$, we have
$\nabla G(\mathbf{W})=\left[\nabla_{\mathbf{w}_{1}}g_{1}(\mathbf{w}_{1}),...,\nabla_{\mathbf{w}_{I}}g_{I}(\mathbf{w}_{I}),\mathbf{0}\right],$ (42)
where $\nabla_{\mathbf{w}_{i}}g_{i}(\mathbf{w}_{i})=\frac{\mathcal{T}^{*}(\mathbf{X}_{i}\mathbf{X}_{i}^{\top})}{N_{i}}-\alpha\mathbf{S}^{\top}\left(\frac{1}{\mathbf{S}\mathbf{w}_{i}}\right)+4\beta\mathbf{w}_{i}$, $\mathcal{T}:\mathbb{R}^{p}\to\mathbb{R}^{d\times d}$ is a linear operator transforming a vector into a Laplacian matrix, and $\mathcal{T}^{*}:\mathbb{R}^{d\times d}\to\mathbb{R}^{p}$ is the adjoint operator of $\mathcal{T}$. See Definition 2 and definition 3 in [49] for more details of $\mathcal{T}$ and $\mathcal{T}^{*}$. Since the regularizer $R(\cdot)$ is decomposable, its dual norm is defined as [32]
$R^{*}(\mathbf{W})=\max\left(\lVert\mathbf{w}_{\mathrm{con}}\rVert_{\infty},\underset{j=1,...,p}{\max}\sqrt{\widetilde{\mathbf{w}}_{j}^{\top}\mathbf{L}_{m}^{-1}\widetilde{\mathbf{w}}_{j}}\right).$ (43)
Based on the equivalence of vector norms, we have
$\underset{j=1,...,p}{\max}\sqrt{\widetilde{\mathbf{w}}_{j}^{\top}\mathbf{L}_{m}^{-1}\widetilde{\mathbf{w}}_{j}}\leq\omega_{\min}(\mathbf{L}_{m})\underset{j=1,...,p}{\max}\lVert\widetilde{\mathbf{w}}_{j}\rVert_{2}$
$\leq$ $\omega_{\min}(\mathbf{L}_{m})\underset{j=1,...,p}{\max}\sqrt{I+1}\lVert\widetilde{\mathbf{w}}_{j}\rVert_{\infty}$
$=$ $\omega_{\min}(\mathbf{L}_{m})\sqrt{I+1}\underset{i=1,...,I+1}{\max}\lVert\mathbf{w}_{i}\rVert_{\infty}.$ (44)
Therefore, we have
$R^{*}(\nabla G(\mathbf{W}^{*}))$
$\leq$ $\omega_{\min}(\mathbf{L}_{m})\sqrt{I+1}\underset{i=1,...,I}{\max}{\left\lVert\nabla_{\mathbf{w}_{i}}G(\mathbf{W}^{*})\right\rVert_{\infty}}$
$=$ $\omega_{\min}(\mathbf{L}_{m})\sqrt{I+1}$
$\underset{i=1,...,I}{\max}\left\lVert\frac{1}{N_{i}}\mathcal{T}^{*}(\mathbf{X}_{i}\mathbf{X}_{i}^{\top})-\alpha\mathbf{S}^{\top}\left(\frac{1}{\mathbf{S}\mathbf{w}^{*}_{i}}\right)+4\beta\mathbf{w}^{*}_{i}\right\rVert_{\infty}$
$\leq$ $\omega_{\min}(\mathbf{L}_{m})\sqrt{I+1}$
$\left(\sqrt{\frac{2d(d-1)}{B_{d}}}+4\beta B_{w^{*}}+\underset{i=1,...,I}{\max}4\left\lVert\frac{\mathbf{X}_{i}\mathbf{X}_{i}^{\top}}{N_{i}}\right\rVert_{\infty}\right).$ (45)
The first inequality of (45) holds due to (44) and $\nabla_{\mathbf{w}_{\mathrm{con}}}G(\mathbf{W}^{*})=\mathbf{0}$. The last inequality holds due to triangle inequality and the definition of $\mathcal{T}^{*}$. Furthermore, according to Lemma 1 in [50], we have
$\mathrm{Pr}\left(\left\lVert\mathbf{\widehat{\Sigma}}_{i}-\varpi(\mathbf{L}_{i}^{*})\right\rVert_{\infty}>\delta\right)$
$\leq$ $4\exp\left(-\frac{N_{i}\delta_{s}^{2}}{2^{7}5^{2}\max_{k}(\varpi^{2}(\mathbf{L}_{i}^{*}){[kk]})}\right),$ (46)
for all $\delta\in(0,40\max_{k}(\varpi^{2}(\mathbf{L}_{i}^{*}){[kk]}))$, where $\widehat{\mathbf{\Sigma}}_{i}=\frac{\mathbf{X}_{i}\mathbf{X}_{i}^{\top}}{N}$. Taking the union bound, for $N_{\min}\geq 2\ln d$, we have
$\max_{i}\lVert\mathbf{\widehat{\Sigma}}_{i}\rVert_{\infty}-\max_{i}\left\lVert\varpi(\mathbf{L}_{i}^{*})\right\rVert_{\infty}\leq 40\sqrt{2}\max_{i,k}\varpi(\mathbf{L}_{i}^{*}){[kk]}\sqrt{\frac{\ln d}{N_{\min}}},$ (47)
with probability at least $1-\frac{4I}{d}$. Bringing (47) back to (45), we complete the proof.
(ii) The function $G(\mathbf{W})$ has the Taylor expansion at $\mathbf{W}^{*}$ as
$G(\mathbf{W}^{*}+\mathbf{\Delta})-G(\mathbf{W}^{*})-\left\langle\nabla G(\mathbf{W}^{*}),\mathbf{\Delta}\right\rangle$
$=$ $\frac{1}{2}\sum_{i=1}^{I}\mathbf{\Delta}_{i}^{\top}\nabla^{2}G(\mathbf{w}_{i}^{*})\mathbf{\Delta}_{i},$ (48)
where $\mathbf{\Delta}_{i}$ is the $i$-th column of $\mathbf{\Delta}$. It is not difficult to check that $\nabla^{2}G(\mathbf{w}_{i}^{*})=4\beta\mathbf{I}+\alpha\mathbf{S}^{\top}\mathrm{diag}\left((\mathbf{S}\mathbf{w}_{i}^{*})^{-2}\right)\mathbf{S}$. Therefore, we have
$G(\mathbf{W}^{*}+\mathbf{\Delta})-G(\mathbf{W}^{*})-\left\langle\nabla G(\mathbf{W}^{*}),\mathbf{\Delta}\right\rangle$
$=$ $\frac{1}{2}\sum_{i=1}^{I}\mathbf{\Delta}_{i}^{\top}\left(4\beta+\alpha\mathbf{S}^{\top}\mathrm{diag}\left((\mathbf{S}\mathbf{w}_{i}^{*})^{-2}\right)\mathbf{S}\right)\mathbf{\Delta}_{i}+0\cdot\mathbf{\Delta}_{I+1}^{\top}\mathbf{\Delta}_{I+1}$
$\geq$ $2\beta\sum_{i=1}^{I+1}\mathbf{\Delta}_{i}^{\top}\mathbf{\Delta}_{i}=2\beta\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}^{2}.$ (49)
Finally, we complete the proof. ∎
Lemma 3.
Define the set $\mathcal{Q}:=\{\mathbf{\Delta}\in\mathbb{R}^{p\times(I+1)}:R(\mathbf{\Delta}_{\mathcal{M}^{\perp}})\leq 3R(\mathbf{\Delta}_{\mathcal{M}})\}$. Suppose $\lambda\geq 2\kappa$, then we have $\widehat{\mathbf{\Delta}}\in\mathcal{Q}$. In addition, let $\epsilon_{d}>0$ be an arbitrary constant. If $\phi(\mathbf{\Delta})>0$ for all elements $\mathbf{\Delta}\in\mathcal{Q}\cap\{\mathbf{\Delta}\in\mathbb{R}^{p\times(I+1)}:\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}=\epsilon_{d}\}$, then $\lVert\widehat{\mathbf{\Delta}}\rVert_{\mathrm{F}}\leq\epsilon_{d}$.
Proof:
The proof follows Lemma 1 and Lemma 4 in [32]. ∎
Now, we apply Lemmas 1-3 to derive the estimation error bound in Theorem 1. First, we need to compute a lower bound for $\phi(\mathbf{\Delta})$. Based on (36) and (41), we have
$\phi(\mathbf{\Delta})$
$\geq$ $\left\langle\nabla G(\mathbf{W}^{*}),\mathbf{\Delta}\right\rangle+\lambda\left(R(\mathbf{\Delta}_{\mathcal{M}^{\perp}})-R(\mathbf{\Delta}_{\mathcal{M}})\right)+2\beta\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}^{2}$
$\geq$ $-\left\lvert\left\langle\nabla G(\mathbf{W}^{*}),\mathbf{\Delta}\right\rangle\right\rvert+\lambda\left(R(\mathbf{\Delta}_{\mathcal{M}^{\perp}})-R(\mathbf{\Delta}_{\mathcal{M}})\right)+2\beta\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}^{2},$ (50)
By the Cauchy-Schwarz inequality, we obtain
$\left\lvert\left\langle\nabla G(\mathbf{W}^{*}),\mathbf{\Delta}\right\rangle\right\rvert\leq R^{*}(\nabla G(\mathbf{W}^{*}))R(\mathbf{\Delta})\leq\kappa R(\mathbf{\Delta}),$ (51)
where the last inequality holds due to Lemma 2 (i). Therefore, if $\lambda\geq 2\kappa$, we can obtain
$\left\lvert\left\langle\nabla G(\mathbf{W}^{*}),\mathbf{\Delta}\right\rangle\right\rvert\leq\frac{\lambda}{2}R(\mathbf{\Delta})=\frac{\lambda}{2}\left(R(\mathbf{\Delta}_{\mathcal{M}^{\perp}})+R(\mathbf{\Delta}_{\mathcal{M}})\right),$ (52)
where the last inequality holds due to the decomposition of the regularizer. Bringing (52) back to (50), we have
$\phi(\mathbf{\Delta})$
$\geq$ $\frac{\lambda}{2}\left(R(\mathbf{\Delta}_{\mathcal{M}^{\perp}})-3R(\mathbf{\Delta}_{\mathcal{M}})\right)+2\beta\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}^{2}$
$\geq$ $-\frac{3\lambda}{2}R(\mathbf{\Delta}_{\mathcal{M}})+2\beta\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}^{2}$
$\geq$ $-\frac{3\lambda}{2}\left(\sqrt{s}+\nu\sqrt{Ip\omega_{\max}(\mathbf{L}_{m})}\right)\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}+2\beta\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}^{2},$ (53)
where the last inequality holds due to (37). Based on Lemma 3, we need to find an $\epsilon_{d}>0$ such that for all $\mathbf{\Delta}\in\mathcal{Q}\cap\{\mathbf{\Delta}\in\mathbb{R}^{p\times(I+1)}:\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}=\epsilon_{d}\}$, $\phi(\mathbf{\Delta})>0$. The right-hand side of (53) is a quadratic form in $\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}$. Thus, we have $\phi(\mathbf{\Delta})>0$ only if
$\lVert\mathbf{\Delta}\rVert_{\mathrm{F}}\geq\frac{3\lambda\left(\sqrt{s}+\nu\sqrt{Ip\omega_{\max}(\mathbf{L}_{m})}\right)}{4\beta}:=\epsilon_{d}.$ (54)
Suppose that $\lambda>2\kappa$ , we complete the proof based on Lemma 2 (i).
Appendix B Proof of Proposition 1
The proof is mainly from [51] but with some modifications. For two vectors $\mathbf{y}$ and $\mathbf{y}^{\prime}$ in $\widetilde{\mathcal{W}}$, we have that
$\left\lVert\nabla_{\mathbf{y}}f^{(t)}_{i}(\mathbf{y},\mathbf{w}_{\mathrm{con}}^{(t)})-\nabla_{\mathbf{y}^{\prime}}f^{(t)}_{i}(\mathbf{y}^{\prime},\mathbf{w}_{\mathrm{con}}^{(t)})\right\rVert_{2}$
$=$ $\left\lVert(4\beta+\rho\gamma_{i}^{(t)})(\mathbf{y}-\mathbf{y}^{\prime})-\alpha\mathbf{S}^{\top}\left(\frac{1}{\mathbf{S\mathbf{y}}+\zeta\mathbf{1}}-\frac{1}{\mathbf{S\mathbf{y}^{\prime}}+\zeta\mathbf{1}}\right)\right\rVert_{2}$
$\leq$ $(4\beta+\rho\gamma_{i}^{(t)})\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{2}+\alpha\left\lVert\mathbf{S}^{\top}\right\rVert_{2}\left\lVert\frac{1}{\mathbf{S\mathbf{y}}+\zeta\mathbf{1}}-\frac{1}{\mathbf{S\mathbf{y}^{\prime}}+\zeta\mathbf{1}}\right\rVert_{2}$
$\leq$ $4\beta\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{2}+\frac{\alpha\lVert\mathbf{S}\rVert_{2}}{\zeta^{2}}\left\lVert\mathbf{S}\mathbf{y}-\mathbf{S}\mathbf{y}^{\prime}\right\rVert_{2}+\rho\gamma_{i}^{(t)}\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{2}$
$\leq$ $4\beta\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{2}+\frac{\alpha\lVert\mathbf{S}\rVert_{2}^{2}}{\zeta^{2}}\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{2}+\rho\gamma_{i}^{(t)}\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{2}$
$=$ $\left(4\beta+\frac{2\alpha(d-1)}{\zeta^{2}}+\rho\gamma_{i}^{(t)}\right)\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert_{2}:=L_{i}^{(t)}\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert_{2},$ (55)
The first inequality holds due to the triangle inequality, while the penultimate equality holds due to Lemma 1 in [51]. From (55), we can conclude that $f_{i}^{(t)}$ is $L_{i}^{(t)}$-Lipschitz smooth w.r.t. $\mathbf{w}_{i}$ on $\widetilde{\mathcal{W}}$. On the other hand, we obtain that
$\left\lVert\nabla_{\mathbf{y}}f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{y})-\nabla_{\mathbf{y}^{\prime}}f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{y}^{\prime})\right\rVert_{2}=\rho\gamma_{i}^{(t)}\left\lVert\mathbf{y}-\mathbf{y}^{\prime}\right\rVert_{2}$ (56)
Therefore, $f_{i}^{(t)}$ is $\rho\gamma_{i}^{(t)}$-Lipschitz smooth w.r.t. $\mathbf{w}_{\mathrm{con}}$.
Appendix C Proof of Theorem 2
For notational simplicity, we fix $K_{i}^{(t)}=1$ and omit the epoch index $k$ in local clients (only one epoch of update is performed in local clients). We remark that the proof can be easily extended to general cases where $K_{i}^{(t)}>1$. Before we start the proof, some lemmas are provided.
Lemma 4.
Suppose that Assumptions 3 and 4 hold, for any client $\mathcal{C}_{i}$, the updated $\mathbf{w}_{i}^{(t+1)}$ satisfies that
$f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t)}_{\mathrm{con}})\leq$ $f_{i}^{(t)}(\mathbf{w}_{i}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})-\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}$
$+\frac{\xi^{2}}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i}^{(0)}\rVert_{2}^{2}$ (57)
Proof:
Recall that $f_{i}^{(t)}(\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}^{(t)})$ is convex and $L_{i}^{(t)}$-Lipschitz smooth w.r.t. $\mathbf{w}_{i}$ as Proposition 1 states. Thus in the $t$-th iteration, given $\gamma_{1}^{(t)},...,\gamma_{I}^{(t)}$, we obtain that
$f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t)}_{\mathrm{con}})\leq f_{i}^{(t)}(\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})+\frac{L_{i}^{(t)}}{2}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\rVert_{2}^{2}$
$+\left\langle\nabla_{\mathbf{w}_{i}}f_{i}^{(t)}(\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}}),\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\right\rangle$ (58)
$f_{i}^{(t)}(\mathbf{w}_{i}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})\geq f_{i}^{(t)}(\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})$
$+\left\langle\nabla_{\mathbf{w}_{i}}f_{i}^{(t)}(\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}}),\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\right\rangle,$ (59)
where the first inequality holds because of the definition of $L_{i}^{(t)}$-Lipschitz smooth, and the second inequality holds since $f_{i}^{(t)}(\mathbf{w}_{i},\mathbf{w}_{\mathrm{con}}^{(t)})$ is a convex function w.r.t $\mathbf{w}_{i}$. Combining the two inequalities, we have
$f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t)}_{\mathrm{con}})$
$\leq$ $f_{i}^{(t)}(\mathbf{w}_{i}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})+\left\langle\nabla_{\mathbf{w}_{i}}f_{i}^{(t)}(\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}}),\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\right\rangle$
$+\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\rVert_{2}^{2}$
$\leq$ $f_{i}^{(t)}(\mathbf{w}_{i}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})+\left\langle\frac{1}{\eta_{w}}(\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}),\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i}^{(t+1)}\right\rangle$
$+\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\rVert_{2}^{2}$
$\leq$ $f_{i}^{(t)}(\mathbf{w}_{i}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})-\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}+\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\rVert_{2}^{2}$
$=$ $f_{i}^{(t)}(\mathbf{w}_{i}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})-\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}$
$+\frac{\xi^{2}}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i}^{(t-1)}\rVert_{2}^{2}$ (60)
where the first inequality holds due to $\eta_{w}<1/L_{max}$. The second inequality holds since updating $\mathbf{w}_{i}^{(t+1)}$ via (23) and (24) is equivalent to the following problem
$\mathbf{w}_{i}^{(t+1)}=$ $\underset{\mathbf{y}\in\mathcal{W}}{\mathrm{argmin}}\,\,\left\langle\nabla_{\mathbf{w}_{i}}{f}_{i}^{(t)}(\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}}),\mathbf{y}\right\rangle$
$+\frac{1}{2\eta_{w}}\lVert\mathbf{y}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\rVert_{2}^{2},$ (61)
which satisfies the following first-order optimality condition for any $\mathbf{y}\in\mathcal{W}$
$\left\langle\nabla_{\mathbf{w}_{i}}{f}_{i}^{(t)}(\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}})+\frac{1}{\eta_{w}}\left(\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\right),\mathbf{y}-\mathbf{w}_{i}^{(t+1)}\right\rangle\geq 0.$ (62)
The third inequality of (60) holds since
$\left\langle\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)},\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i}^{(t+1)}\right\rangle$
$=$ $\frac{1}{2}\left(\lVert\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\rVert_{2}^{2}-\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i,\mathrm{ex}}^{(t)}\rVert_{2}^{2}-\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}\right).$ (63)
Finally, we complete the proof. ∎
Lemma 5.
Suppose Assumptions 3-4 hold, the updated $\mathbf{w}^{(t+1)}_{\mathrm{con}}$ satisfies that
$\sum_{i}^{I}f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t+1)}_{\mathrm{con}})+\lambda\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}\rVert_{1}$
$\leq$ $\sum_{i}^{I}f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t)}_{\mathrm{con}})+\lambda\lVert\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{1}-\frac{\rho C_{\gamma}^{(t)}}{2}\left\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\right\rVert_{2}^{2}.$ (64)
Proof:
For simplicity, we first define $F(\mathbf{w}_{\mathrm{con}})=\sum_{i}^{I}f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}_{\mathrm{con}})$, and $\nabla_{\mathbf{w}_{\mathrm{con}}}F(\mathbf{w}^{(t)}_{\mathrm{con}})$ is simplified as $\nabla F(\mathbf{w}^{(t)}_{\mathrm{con}})$. Since $f_{i}^{(t)}(\mathbf{w}_{i}^{(t+1)},\mathbf{w}_{\mathrm{con}})$ is $\rho\gamma_{i}^{(t)}$-convex w.r.t. $\mathbf{w}_{\mathrm{con}}$, $F(\mathbf{w}_{\mathrm{con}})$ is $\rho C_{\gamma}^{(t)}$ convex w.r.t. $\mathbf{w}_{\mathrm{con}}$, and we obtain
$F(\mathbf{w}^{(t+1)}_{\mathrm{con}})\leq$ $F(\mathbf{w}^{(t)}_{\mathrm{con}})+\left\langle\nabla F(\mathbf{w}^{(t)}_{\mathrm{con}}),\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\right\rangle$
$+\frac{\rho C_{\gamma}^{(t)}}{2}\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{2}^{2}.$ (65)
On the other hand, due to the convexity of $F(\mathbf{w}^{(t+1)}_{\mathrm{con}})$, for any $\mathbf{y}\in\mathcal{W}$, we have $F(\mathbf{y})\geq F(\mathbf{w}^{(t)}_{\mathrm{con}})+\left\langle\nabla F(\mathbf{w}^{(t)}_{\mathrm{con}}),\mathbf{y}-\mathbf{w}^{(t)}_{\mathrm{con}}\right\rangle$. Combining it with (65), we finally reach that
$F(\mathbf{w}^{(t+1)}_{\mathrm{con}})\leq$ $F(\mathbf{y})-\left\langle\nabla F(\mathbf{w}^{(t)}_{\mathrm{con}}),\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{y}\right\rangle$
$+\frac{\rho C_{\gamma}^{(t)}}{2}\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{2}^{2}.$ (66)
We next focus on $h(\mathbf{w}_{\mathrm{con}})=\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}$. Recall that $\mathbf{w}^{(t+1)}_{\mathrm{con}}=\mathrm{prox}_{\mu^{(t)}\lVert\cdot\rVert_{1}}\left(\frac{\sum_{i=1}^{I}\gamma_{i}^{(t)}\mathbf{w}_{i}^{(t+1)}}{C_{\gamma}^{(t)}}\right)={\mathrm{argmin}}_{\mathbf{w}_{\mathrm{con}}}\,\frac{1}{2}\left\lVert\mathbf{w}_{\mathrm{con}}-\frac{\sum_{i=1}^{I}\gamma_{i}^{(t)}\mathbf{w}_{i}^{(t+1)}}{C_{\gamma}^{(t)}}\right\rVert_{2}^{2}+\frac{\lambda}{\rho C_{\gamma}^{(t)}}\lVert\mathbf{w}_{\mathrm{con}}\rVert_{1}$. Define $\frac{\sum_{i=1}^{I}\gamma_{i}^{(t)}\mathbf{w}_{i}^{(t+1)}}{C_{\gamma}^{(t)}}={\mathbf{w}_{\mathrm{avg}}^{(t+1)}}$, by optimal conditions we obtain
$\mathbf{0}\in\frac{\rho{C_{\gamma}^{(t)}}}{\lambda}\left(\mathbf{w}^{(t+1)}_{\mathrm{con}}-{\mathbf{w}_{\mathrm{avg}}^{(t+1)}}\right)+\partial h(\mathbf{w}^{(t+1)}_{\mathrm{con}})$
$\implies$ $\frac{\rho{C_{\gamma}^{(t)}}}{\lambda}\left({\mathbf{w}_{\mathrm{avg}}^{(t+1)}}-\mathbf{w}^{(t+1)}_{\mathrm{con}}\right)\in\partial h(\mathbf{w}^{(t+1)}_{\mathrm{con}})$ (67)
Similarly, due to the convexity of $h(\mathbf{w}_{\mathrm{con}})$, we have $h(\mathbf{y})\geq h(\mathbf{w}^{(t+1)}_{\mathrm{con}})+\left\langle\partial h(\mathbf{w}^{(t+1)}_{\mathrm{con}}),\mathbf{y}-\mathbf{w}^{(t+1)}_{\mathrm{con}}\right\rangle$. Combining it with (66) and (67), we obtain
$F(\mathbf{w}^{(t+1)}_{\mathrm{con}})+\lambda h(\mathbf{w}^{(t+1)}_{\mathrm{con}})\leq F(\mathbf{y})+\lambda h(\mathbf{y})+\big{\langle}\nabla F(\mathbf{w}^{(t)}_{\mathrm{con}})$
$+{\rho}{C_{\gamma}^{(t)}}\left({\mathbf{w}_{\mathrm{avg}}^{(t+1)}}-\mathbf{w}^{(t+1)}_{\mathrm{con}}\right),\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{y}\big{\rangle}$
$+\frac{\rho C_{\gamma}^{(t)}}{2}\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{2}^{2}$
$\implies$ $F(\mathbf{w}^{(t+1)}_{\mathrm{con}})+\lambda h(\mathbf{w}^{(t+1)}_{\mathrm{con}})\leq F(\mathbf{y})+\lambda h(\mathbf{y})$
$+\rho{C_{\gamma}^{(t)}}\left\langle\mathbf{w}^{(t)}_{\mathrm{con}}-\mathbf{w}^{(t+1)}_{\mathrm{con}},\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{y}\right\rangle$
$+\frac{\rho C_{\gamma}^{(t)}}{2}\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{2}^{2}$
$\stackrel{{\scriptstyle\mathbf{y}=\mathbf{w}^{(t)}_{\mathrm{con}}}}{{\Longrightarrow}}$ $F(\mathbf{w}^{(t+1)}_{\mathrm{con}})+\lambda h(\mathbf{w}^{(t+1)}_{\mathrm{con}})\leq F(\mathbf{w}^{(t)}_{\mathrm{con}})+\lambda h(\mathbf{w}^{(t)}_{\mathrm{con}})$
$-\frac{\rho C_{\gamma}^{(t)}}{2}\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}-\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{2}^{2}$ (68)
Finally, we complete the proof. ∎
Equipped with the two lemmas, we start our proof of Theorem 2. We define some symbols to simplify the expression. Specifically,
$E^{(t)}_{l}=$ $\sum_{i=1}^{I}f_{i}^{(t)}\left(\mathbf{w}_{i}^{(t)},\mathbf{w}^{(t)}_{\mathrm{con}}\right)-f_{i}^{(t)}\left(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t)}_{\mathrm{con}}\right)$
$E^{(t)}_{c}=$ $\sum_{i=1}^{I}f_{i}^{(t)}\left(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t)}_{\mathrm{con}}\right)+\lambda\lVert\mathbf{w}^{(t)}_{\mathrm{con}}\rVert_{1}$
-- $\sum_{i=1}^{I}f_{i}^{(t)}\left(\mathbf{w}_{i}^{(t+1)},\mathbf{w}^{(t+1)}_{\mathrm{con}}\right)-\lambda\lVert\mathbf{w}^{(t+1)}_{\mathrm{con}}\rVert_{1}$
$e^{(t)}_{l}=$ $\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}-\frac{\xi^{2}}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i}^{(t-1)}\rVert_{2}^{2}$
$e^{(t)}_{c}=$ $\frac{\rho C_{\gamma}^{(t)}}{2}\lVert\mathbf{w}^{(t)}_{\mathrm{con}}-\mathbf{w}^{(t+1)}_{\mathrm{con}}\rVert_{2}^{2}$ (69)
Then combining Lemma 4 and Lemma 5, we have
$\sum_{i=1}^{I}e^{(t)}_{l}+e^{(t)}_{c}\leq E^{(t)}_{l}+E^{(t)}_{c}.$ (70)
Summing (70) from 0 to $T-1$, we obtain
$\sum_{t=0}^{T-1}\sum_{i=1}^{I}e^{(t)}_{l}+e^{(t)}_{c}\leq\sum_{t=0}^{T-1}E^{(t)}_{l}+E^{(t)}_{c}$
$=$ $\sum_{i=1}^{I}f_{i}^{(0)}\left(\mathbf{w}_{i}^{(0)},\mathbf{w}^{(0)}_{\mathrm{con}}\right)-f_{i}i^{(T)}\left(\mathbf{w}_{i}^{(T)},\mathbf{w}^{(T)}_{\mathrm{con}}\right)$
$+\lambda\lVert\mathbf{w}_{\mathrm{con}}^{(0)}\rVert_{1}-\lambda\lVert\mathbf{w}_{\mathrm{con}}^{(T)}\rVert_{1}$
$\leq$ $\sum_{i=1}^{I}f_{i}^{(0)}\left(\mathbf{w}_{i}^{(0)},\mathbf{w}^{(0)}_{\mathrm{con}}\right)-f_{i}^{*}+\lambda\lVert\mathbf{w}_{\mathrm{con}}^{(0)}-\mathbf{w}_{\mathrm{con}}^{(T)}\rVert_{1}$
$\leq$ $\sum_{i=1}^{I}f_{i}^{(0)}\left(\mathbf{w}_{i}^{(0)},\mathbf{w}^{(0)}_{\mathrm{con}}\right)-f_{i}^{*}+\lambda\sqrt{p}\lVert\mathbf{w}_{\mathrm{con}}^{(0)}-\mathbf{w}_{\mathrm{con}}^{(T)}\rVert_{2}$
$\leq$ $\sum_{i=1}^{I}f_{i}^{(0)}\left(\mathbf{w}_{i}^{(0)},\mathbf{w}^{(0)}_{\mathrm{con}}\right)-f_{i}^{*}+\lambda\sqrt{p}B_{r}\triangleq\Upsilon.$ (71)
Notice that
$\sum_{t=0}^{T-1}e^{(t)}_{l}$
$=$ $\sum_{t=0}^{T-1}\left(\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}-\frac{\xi^{2}}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t)}-\mathbf{w}_{i}^{(t-1)}\rVert_{2}^{2}\right)$
$\geq$ $\sum_{t=0}^{T-1}\frac{1-\xi^{2}}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}-\frac{1}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(0)}-\mathbf{w}_{i}^{(-1)}\rVert_{2}^{2}$
$=$ $\sum_{t=0}^{T-1}\frac{1-\xi^{2}}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}.$ (72)
We expand the summation operator to obtain the first inequality. The last equality holds due to initialization conditions of our algorithm. With (71) and (72), we obtain that
$\frac{1}{T}\sum_{t=0}^{T-1}\sum_{i=1}^{I}\frac{1-\xi^{2}}{2\eta_{w}}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}+\frac{1}{T}\sum_{t=0}^{T-1}e^{(t)}_{c}\leq\frac{\Upsilon}{T},$ (73)
Since $10\leq\xi<$ 1, we can conclude that
$\frac{1}{T}\sum_{t=0}^{T-1}\left(\sum_{i=1}^{I}\lVert\mathbf{w}_{i}^{(t+1)}-\mathbf{w}_{i}^{(t)}\rVert_{2}^{2}\right)=\frac{C_{1}}{T},$ (74)
where $C_{1}=\frac{2\Upsilon\eta_{w}}{1-\xi^{2}}$.
In the same way, we obtain that
$\frac{1}{T}\sum_{t=0}^{T-1}e^{(t)}_{c}\leq\frac{\Upsilon}{T}$
$\implies$ $\frac{\rho}{2T}\sum_{t=0}^{T-1}C_{\gamma}^{(t)}\lVert\mathbf{w}^{(t)}_{\mathrm{con}}-\mathbf{w}^{(t+1)}_{\mathrm{con}}\rVert_{2}^{2}\leq\frac{\Upsilon}{T}$
$\implies$ $\frac{\rho I\gamma_{\min}}{2T}\sum_{t=0}^{T-1}\lVert\mathbf{w}^{(t)}_{\mathrm{con}}-\mathbf{w}^{(t+1)}_{\mathrm{con}}\rVert_{2}^{2}\leq\frac{\Upsilon}{T}$
$\implies$ $\frac{1}{T}\sum_{t=0}^{T-1}\lVert\mathbf{w}^{(t)}_{\mathrm{con}}-\mathbf{w}^{(t+1)}_{\mathrm{con}}\rVert_{2}^{2}\leq\frac{C_{2}}{T}$ (75)
where $C_{2}=\frac{2\Upsilon}{\rho I\gamma_{\min}}$. The third inequality holds since $\gamma^{(t)}_{i}\geq\gamma_{\min}$ for all $i$ and $t$.